solve the following equation by changing the independent variable x^2 d^2 y/dx - dy/dx -4x^3 y = 8x^3 sin x^2 , x>0
1
Expert's answer
2020-06-08T20:46:40-0400
"x{d^2y\over dx^2}-{dy\over dx}-4x^3y=8x^3\sin x^2,x>0"

"{d^2y\over dx^2}-{1\over x}\cdot{dy\over dx}-4x^2y=8x^2\sin x^2"

Here

"P=-\dfrac{1}{x},"


"Q=-4x^2,"


"R=8x^2\sin x^2"

Changing the independent variable from "x" to "z"

"{d^2y\over dz^2}+P_1{dy\over dz}+Q_1y=R_1"

where


"P_1=\dfrac{{d^2z\over dx^2}+P{dz\over dx}}{({dz\over dx})^2}"

"Q_1=\dfrac{Q}{({dz\over dx})^2}"

"R_1=\dfrac{R}{({dz\over dx})^2}"

We choose "z" such that


"Q_1=\dfrac{Q}{({dz\over dx})^2}=constant"

Let


"Q_1=\dfrac{-4x^2}{({dz\over dx})^2}=-1"

Then


"{dz\over dx}=2x"

integrating


"z=x^2"

We have


"{d^2z\over dx^2}=2"

"P_1=\dfrac{{d^2z\over dx^2}+P{dz\over dx}}{({dz\over dx})^2}=\dfrac{2+(-\dfrac{1}{x})(2x)}{(2)^2}=0"

"R_1=\dfrac{R}{({dz\over dx})^2}=\dfrac{8x^2\sin x^2}{(2x)^2}=2\sin z"

"{d^2y\over dz^2}-y=2\sin z"

Method of undetermined coefficients

Homogeneous second order differentional equation


"{d^2y\over dz^2}-y=0"

Characteristic equation


"\lambda^2-1=0""\lambda_1=-1,\lambda_2=1"

The general solution of the homogeneous differential equation is


"y_0=C_1e^{-z}+C_2e^{z}"

Let "Y=A\cos z+B\sin z," then


"{dY\over dz}=-A\sin z+B\cos z,"

"{d^2Y\over dz^2}=-A\cos z-B\sin z"

"A\cos z+B\sin z+A\cos z+B\sin z=2\sin z"

"A=0"

"B=1"


"Y=\sin z"

The general solution of the nonhomogeneous equation


"y(z)=y_0+Y"

"y(z)=C_1e^{-z}+C_2e^{z}+\sin z"

Substitute "z=x^2" and obtain the general solution


"y(z)=C_1e^{-x^2}+C_2e^{x^2}+\sin x^2"